OpenStudy (anonymous):

3 x - 9 x + 7 - 3 = -8 solve for x, polpak

OpenStudy (anonymous):

-6x=-12 x=2

OpenStudy (anonymous):

Start by adding 3 to -8. Then subtract 7 from the left side of the equation and then the right.. sooo.... 3x-9x+7=-5 3x-9x=-12 Then factor out a three from the x.. 3(x-3x)=-12

OpenStudy (anonymous):

Ankur, stop giving people solutions they could just get directly from wolframalpha, etc. They want help, not answers. Nobody cares if you know how to do it.

OpenStudy (anonymous):

Then you divide by three to get rid of of the product.. x-3x=-4

OpenStudy (anonymous):

@ankur504Yes, Please stop....I find it aggravating...

OpenStudy (anonymous):

Are you following Artist?

OpenStudy (anonymous):

why cant u stop teaching other people.....I WILL DO WHATEVER I WANT.....PERIOD

OpenStudy (anonymous):

@ Kandyse, Kinda...

OpenStudy (anonymous):

I don't think you need to factor out a 3 so much as you just need to combine your x's

OpenStudy (anonymous):

Artist, recall that anything you do to one side of an equation you can also do to the other.

OpenStudy (anonymous):

Yeah, then factor out an x, so you have x(1-3)

OpenStudy (anonymous):

then you have -2x

OpenStudy (anonymous):

@ polpak, yes...That is what i was taught...

OpenStudy (anonymous):

then you simply divide -2 to both sides of the equation, because as polpak said, you do it to both sides of the equation :)

OpenStudy (anonymous):

Sooo, then you have x=2 :)

OpenStudy (anonymous):

Thanks! Really helpful!

OpenStudy (anonymous):

:D you are welcome!! :) Glad to help :)

OpenStudy (anonymous):

Good work! =)

OpenStudy (anonymous):

Thanks! I fanned you Kandyse! Could you fan me? Thanks!

OpenStudy (anonymous):

lol, I already have :)

OpenStudy (anonymous):

cool! Thanks!